# Variational Bayesian Bow tie Neural Networks with Shrinkage

### v2

@@ 4.3 UCI regression datasets
Lastly, we consider publicly available datasets from the UCI Machine Learning Repository M. et al. (2007): Boston housing Harrison and Rubinfeld (1978), Energy Tsanas and Xifara (2012), Yacht dynamics J. et al. (2013), Concrete compressive strength Yeh (2007) and Concrete slump test Yeh (2009) (see Section C.4 for the description of the datasets). For all of the UCI regression tasks, we fit a neural network with one hidden layer and $D_{H}=50$ hidden units. Table 3 compares RMSE, NLL and empirical coverage of the observations, for VBNN, SVI and BBB baselines. Our model is more consistent in performance across various datasets than the SVI baseline and is comparable to BBB. In terms of the coverage and uncertainty quantification, VBNN performs consistently well across all datasets in contrast to the other models.
Table 3: RMSE, NLL and Coverage for UCI datasets.
Dataset RMSE NLL Coverage
VBNN SVI BBB VBNN SVI BBB VBNN SVI BBB
Slump $6.99\pm 1.4$ $7.02\pm 1.0$ $\mathbf{6.83\pm 1.52}$ $\mathbf{3.37\pm.2}$ $3.97\pm.6$ $4.33\pm 1.3$$\mathbf{.92\pm.06}$$75\pm.1$$84\pm.14$
Yacht $1.34\pm.3$ $1.73\pm.44$ $\mathbf{1.02\pm.32}$ $1.74\pm.19$ $2.08\pm.12$$\mathbf{.9\pm.49}$$\mathbf{.96\pm.03}$$9$02 .99±.02plus-or-minus.99.02.$9$02
Boston $\mathbf{3.17\pm.5}$ $3.25\pm.54$ $3.18\pm 1.13$ $\mathbf{2.59\pm.14}$ $2.95\pm.4$ $2.69\pm.42$$\mathbf{.96\pm.03}$$82\pm.04$$87\pm.04$
Energy $\mathbf{1.29\pm.19}$ $2.35\pm.26$ 1.88±.33plus-or-minus1.88.331.$8$33 $1.7\pm.14$ $2.31\pm.14$ $\mathbf{1.43\pm.19}$$\mathbf{.98\pm.01}$$89\pm.03$$9$01
Concrete $6$45 $6.04\pm.66$ $\mathbf{5.7\pm.71}$ $3.31\pm.07$ 3.22±.09plus-or-minus3.22.093.$2$09 $\mathbf{3.09\pm 0.22}$$\mathbf{.96\pm.02}$$97\pm.02$$94\pm.03$
@@ 5 Discussion
In this paper, we presented a variational bow tie neural network (VBNN) that is amendable to Polya-gamma data augmentation so that the variational inference can be performed via the CAVI algorithm. While the idea of the stochastic relaxation described in Section 2.1 was introduced in Smith et al. (2021), the novelty of our model is in the employment of the variational inference techniques as well as sparsity-inducing priors. Namely, we implement continuous global-local shrinkage priors and propose a post-process technique for node selection. Additionally, we consider an improvement of the classical CAVI algorithm by adding EM steps for critical hyperparameters. In this way, we enrich the class of models which are handled within the structured mean-field paradigm. We provide all the necessary (computations), techniques, and illustrative experiments demonstrating the utility of the model.
At each iteration, CAVI has to cycle through the entire data set, which can be computationally expensive and inefficient for large sample sizes. An alternative to coordinate ascent is gradient-based optimization and a future direction of this research will extend the algorithm by employing Stochastic variational inference Hoffman et al. (2013) and subsampling. Moreover, an extension to other output types, such as classification tasks, will be developed through additional Polya-gamma augmentation techniques Durante and Rigon (2019).
@@ Appendix A Derivations of the Variational Posterior
×$\times\prod_{l}^{L}\exp\left(-\frac{1}{2}\mathbb{E}\left[\sum_{n}^{N}\sum_{d}^{D_{l}}(\eta_{l,d})^{-2}\left(a_{n,l,d}-\gamma_{n,l,d}\odot z_{n,l,d}\right)^{2}\right]\right)\times\prod_{l}^{L}\prod_{d}^{D_{l}}\left((\eta_{l,d}^{2})^{-\alpha^{h}_{0}-1-\frac{N}{2}}\exp\left(-\frac{\beta^{h}_{0}}{\eta^{2}_{l,d}}\right)\right)$
@@
×$\times\prod_{l}^{L}\prod_{d}^{D_{l}}(\eta_{l,d}^{2})^{-\alpha^{h}_{0}-1-\frac{N}{2}}\exp\left(-\frac{1}{\eta^{2}_{l,d}}\left(\beta^{h}_{0}+\frac{1}{2}\sum_{n}^{N}\mathbb{E}\left[\left(a_{n,l,d}-\gamma_{n,l,d}\odot z_{n,l,d}\right)^{2}\right]\right)\right).$
@@
$\beta_{l,d}$ $=\beta^{h}_{0}+\frac{1}{2}\sum_{n}^{N}\mathbb{E}\left[\left(a_{n,l,d}-\gamma_{n,l,d}\odot z_{n,l,d}\right)^{2}\right],\quad d=1,\ldots,D_{l},\quad l=1,\ldots,L,$
@@
$\mathbb{E}\left[\left(a_{n,l,d}-\gamma_{n,l,d}\cdot z_{n,l,d}\right)^{2}\right]=\sum_{n}^{N}\left(\mathbb{E}\left[a_{n,l,d}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[b_{l,d}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[\mathbf{W}_{l,d}\right]\mathbb{E}\left[\mathbf{a}_{n,l-1}\right]\right)^{2}$
$+\sum_{n}^{N}\mathbb{E}\left[a^{2}_{n,l,d}\right]-\mathbb{E}\left[a_{n,l,d}\right]^{2}+\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[b_{l,d}^{2}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]^{2}\mathbb{E}\left[b_{l,d}\right]^{2}$
@@
$\propto\prod_{n}^{N}\exp\left(\mathbb{E}\left[\log\frac{1}{\sqrt{|\bm{\Sigma}_{L+1}|}}\right]\right)\exp\left(\mathbb{E}\left[-\frac{1}{2}\left(\mathbf{y}_{n}-\mathbf{W}_{L+1}\mathbf{a}_{n,L}-\mathbf{b}_{L+1}\right)^{T}\bm{\Sigma}_{L+1}^{-1}\left(\mathbf{y}_{n}-\mathbf{W}_{L+1}\mathbf{a}_{n,L}-\mathbf{b}_{L+1}\right)\right]\right)$
×$\times\prod_{l}^{L}\prod_{n}^{N}\exp\left(\mathbb{E}\left[\log\frac{1}{\sqrt{|\bm{\Sigma}_{l}|}}\right]\right)\exp\left(\mathbb{E}\left[\left(-\frac{1}{2}\left(\mathbf{a}_{n,l}-\bm{\gamma}_{n,l}\mathbf{W}_{l}\mathbf{a}_{n,l-1}-\bm{\gamma}_{n,l}\mathbf{b}_{l}\right)^{T}\bm{\Sigma}_{l}^{-1}\left(\mathbf{a}_{n,l}-\bm{\gamma}_{n,l}\mathbf{W}_{l}\mathbf{a}_{n,l-1}-\bm{\gamma}_{n,l}\mathbf{b}_{l}\right)\right)\right]\right)$
@@
$\propto\exp\left(-\frac{1}{2}\widetilde{\mathbf{W}}_{L+1,d}\mathbf{D}^{-1}_{L+1,d}\widetilde{\mathbf{W}}_{L+1,d}^{T}-\frac{\mathbb{E}\left[\eta_{L+1,d}^{-2}\right]}{2}\left(\widetilde{\mathbf{W}}_{L+1,d}\left(\sum_{n}^{N}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\widetilde{\mathbf{a}}_{n,L}^{T}\right]\right)\widetilde{\mathbf{W}}_{L+1,d}^{T}-2\widetilde{\mathbf{W}}_{L+1,d}\left(\sum_{n}^{N}y_{n}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\right]\right)\right)\right)$
$\propto\exp\left(-\frac{1}{2}\left(\widetilde{\mathbf{W}}_{L+1,d}\left(\mathbf{D}^{-1}_{L+1,d}+\mathbb{E}\left[\eta_{L+1,d}^{-2}\right]\sum_{n}^{N}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\widetilde{\mathbf{a}}_{n,L}^{T}\right]\right)\widetilde{\mathbf{W}}_{L+1,d}^{T}-2\mathbb{E}\left[\eta_{L+1,d}^{-2}\right]\widetilde{\mathbf{W}}_{L+1,d}\left(\sum_{n}^{N}y_{n}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\right]\right)\right)\right)$
@@
$\left.-\frac{\mathbb{E}\left[\eta_{l,d}^{-2}\right]}{2}\widetilde{\mathbf{W}}_{l,d}\left(\sum_{n}^{N}\mathbb{E}\left[\gamma_{n,l,d}^{2}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)\widetilde{\mathbf{W}}_{l,d}^{T}+\mathbb{E}\left[\eta_{l,d}^{-2}\right]\widetilde{\mathbf{W}}_{l,d}\left(\sum_{n}^{N}\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[a_{n,l,d}\mathbf{a}_{n,l-1}\right]\right)\right.$
@@
$\mathbf{m}_{l,d}^{T}$ $=\mathbf{B}_{l,d}\left(\sum_{n}^{N}\left(\mathbb{E}\left[(\eta_{l,d})^{-2}\right]\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[a_{n,l,d}\widetilde{\mathbf{a}}_{n,l-1}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\left(\mathbb{E}\left[\gamma_{n,l,d}\right]-\frac{1}{2}\right)\right)\right).$
@@
$\propto\prod_{n}^{N}\prod_{l}^{L}\prod_{l}^{D_{l}}\exp\left(-\frac{1}{2\eta_{l,d}^{2}}\mathbb{E}\left[\left(a_{n,l,d}-\gamma_{n,l,d}\left(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d}\right)\right)^{2}\right]+\mathbb{E}\left[\frac{\gamma_{n,l,d}(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d})}{T}\right]\right).$
@@
$q(\gamma_{n,l,d})\propto\exp\left(-\frac{1}{2}\mathbb{E}\left[\eta_{l,d}^{-2}\right]\left(\gamma_{n,l,d}^{2}\mathbb{E}\left[\left(\widetilde{\mathbf{W}}_{l,d}\widetilde{\mathbf{a}}_{n,l-1}\right)^{2}\right]-2\gamma_{n,l,d}\mathbb{E}\left[a_{n,l,d}\widetilde{\mathbf{W}}_{l,d}\widetilde{\mathbf{a}}_{n,l-1}\right]\right)+\frac{1}{T}\gamma_{n,l,d}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\widetilde{\mathbf{a}}_{n,l-1}\right]\right)$
$\propto\exp\left(\gamma_{n,l,d}\left(-\frac{\mathbb{E}\left[\eta_{l,d}^{-2}\right]}{2}\operatorname{Tr}\left(\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}^{T}\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)+\mathbb{E}\left[\eta_{l,d}^{-2}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}a_{n,l,d}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right)\right)$
@@
$\rho_{n,l,d}=$ $\sigma\left(\mathbb{E}\left[\eta_{l,d}^{-2}\right]\left(-\frac{1}{2}\operatorname{Tr}\left(\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}^{T}\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)+\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}a_{n,l,d}\right]\right)+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right).$
@@
×$\times\prod_{n}^{N}\exp\left(-\frac{1}{2}\sum_{l}^{L}\sum_{d}^{D_{l}}\mathbb{E}\left[\frac{1}{\eta_{l,d}^{2}}\right]\mathbb{E}\left[\left(a_{n,l,d}-\gamma_{n,l,d}\left(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d}\right)\right)^{2}\right]\right)$

### v3

@@ 4.3 UCI Regression Datasets
Figure 10: RMSE (normalized w.r.t. to the standard deviation of the target), NLL and empirical coverage for UCI datasets. When illustrating the coverage, the dashed red line depicts the ideal scenario with empirical coverage equal to $95\%$ CI level.
Lastly, we consider publicly available datasets from the UCI Machine Learning Repository (M. et al., 2007): Boston housing (Harrison and Rubinfeld, 1978), Energy (Tsanas and Xifara, 2012), Yacht dynamics (J. et al., 2013), Concrete compressive strength (Yeh, 2007) and Concrete slump test (Yeh, 2009) (see Section D.4 for the description of the datasets). For all of the UCI regression tasks, we fit a neural network with one hidden layer and $D_{H}=50$ hidden units. Figure 10 compares RMSE, NLL and empirical coverage of the observations across the methods (see also Table 4 in Section D.5), where we additionally consider ensembles of four variational approximations with VBNN and SVBNN (denoted as 4VBNN and 4SVBNN, respectively). Overall, HMC outperforms all the considered methods, but at a much higher cost. VBNN provides an improvement compared to GVBNN, further motivating the choice of sparsity-inducing priors, and while SVBNN offers considerable computational savings, the quality of the approximation deteriorates compared to VBNN. Overall, VBNN has consistently well-calibrated uncertainty quantification and empirical coverage for the observations compared with other variational approaches, and ensembles of VBNNs are competitive with other approaches in terms of the RMSE and NLL.
Figure 11: Slump dataset. Performance in terms of the RMSE, NLL and EC for single models (plain colored) and ensembles (color with hatches) obtained from four parallel runs. RMSE and NLL are scaled with respect to the best model (top row). The relative performance (bottom right) is illustrated on the log-scale, and color reflects if ensembles improved the metric (i.e. bar with hatches illustrates the scale of improvements obtained with ensembles, conversely, bar without the hatches illustrates the scale at which single run outperformed ensembles).
Further, we consider slump dataset and implement ensembles of 4 parallel runs of all of the considered methods (4mfVI, 4BBB, 4HSBNN, 4GVBNN, 4VBNN, 4SVBNN) and compare the results to approximations obtained in a single run (mfVI, BBB, HSBNN, GVBNN, VBNN, SVBNN). We do not consider HMC in this experiment due to its computational costs. Figure 11 compares RMSE, NLL, EC of the 12 methods and additionally illustrates relative differences among approaches, where for RMSE and NLL, we consider absolute relative differences between ensembles and single runs, and for empirical coverage of the observations, we illustrate the absolute deviation from the $95\%$ CI. Overall, ensembles improve uncertainty quantification, and in most cases also RMSE and NLL (the only exception being NLL for BBB). While BBB and HSBNN have the lowest RMSE (although with high variability), the NLL and empirical coverage suggest overconfidence, even with ensembles. In contrast, VBNN has an improved balance between accuracy and uncertainty quantification, which is further enhaced by ensembles.
@@ 5 Discussion
In this paper, we presented a variational bow tie neural network (VBNN) that is amendable to Polya-gamma data augmentation so that the variational inference can be performed via the CAVI algorithm. While the idea of the stochastic relaxation described in Section 2.1 was introduced in (Smith et al., 2021), the novelty of our model is in the employment of the variational inference techniques as well as sparsity-inducing priors. Namely, we implement continuous global-local shrinkage priors and propose a post-process technique for node selection. Additionally, we consider an improvement of the classical CAVI algorithm by adding EM steps for critical hyperparameters. In this way, we enrich the class of models which are handled within the structured mean-field paradigm. We provide all the necessary computations, techniques, and illustrative experiments demonstrating the utility of the model. Addressing the scalability with respect to the number of data points, we extend the CAVI algorithm to SVI (Hoffman et al., 2013) which benefits from exploiting natural gradients and subsampling. In the future, we could improve the algorithm by employing an adaptive learning rate which is based on realisations of a noisy estimate of the natural gradient of ELBO with respect to global variational parameters and moving averages (Ranganath et al., 2013; Schaul et al., 2013). Alternatively, instead of changing the learning rate, one could adapt the mini-batch size based on the estimated gradient noise covariance and the magnitude of the gradient (Balles et al., 2017). To address scalability with respect to the network’s width, future work will explore incorporating node selection within the CAVI algorithm when training.
The variational bow tie neural network is also amenable to other prior choices. For example, horseshoe priors can be implemented through the introduction of auxiliary variables to replace each half-Cauchy random variable with the hierarchical formulation based on Inverse-Gamma variables (Wand et al., 2011; Louizos et al., 2017; Ghosh et al., 2018). Additionally, the regularized version of horseshoe priors ("ponyshoe") could be considered, which is known to perform better than the classical horseshoe, especially when the larger coefficients are weakly identified by the data (Piironen and Vehtari, 2017b, a; Ghosh et al., 2019). Finally, an extension to other output types, such as classification tasks, can be developed through additional Polya-gamma augmentation techniques (Durante and Rigon, 2019).
@@ Appendix A Derivations of the Variational Posterior
Assuming hidden layers of the dimension $D$, the computational complexity of updating the global shrinkage variable is $\mathcal{O}(LD\max(D_{0},D,D_{L+1}))$.
@@
Similarly, given hidden layers of the dimension $D$, the computational complexity of updating the global shrinkage variable is $\mathcal{O}(LD\max(D_{0},D,D_{L+1}))$.
@@
×$\times\prod_{l}^{L}\exp\left(-\frac{1}{2}\mathbb{E}\left[\sum_{n}^{N}\sum_{d}^{D_{l}}(\eta_{l,d})^{-2}\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\odot z_{n,l,d}\right)^{2}\right]\right)\times\prod_{l}^{L}\prod_{d}^{D_{l}}\left((\eta_{l,d}^{2})^{-\alpha^{h}_{0}-1-\frac{N}{2}}\exp\left(-\frac{\beta^{h}_{0}}{\eta^{2}_{l,d}}\right)\right)$
@@
×$\times\prod_{l}^{L}\prod_{d}^{D_{l}}(\eta_{l,d}^{2})^{-\alpha^{h}_{0}-1-\frac{N}{2}}\exp\left(-\frac{1}{\eta^{2}_{l,d}}\left(\beta^{h}_{0}+\frac{1}{2}\sum_{n}^{N}\mathbb{E}\left[\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\odot z_{n,l,d}\right)^{2}\right]\right)\right).$
@@
$\beta_{l,d}$ $=\beta^{h}_{0}+\frac{1}{2}\sum_{n}^{N}\mathbb{E}\left[\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\odot z_{n,l,d}\right)^{2}\right],\quad d=1,\ldots,D_{l},\quad l=1,\ldots,L,$
@@
$\mathbb{E}\left[\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\cdot z_{n,l,d}\right)^{2}\right]$ $=\sum_{n}^{N}\left(\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[b_{l,d}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[\mathbf{W}_{l,d}\right]\mathbb{E}\left[\mathbf{a}_{n,l-1}\right]\right)^{2}$
$+\sum_{n}^{N}\mathbb{E}\left[a^{2}_{n,l,d}\right]-\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]^{2}+\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[b_{l,d}^{2}\right]-\mathbb{E}\left[\gamma_{n,l,d}\right]^{2}\mathbb{E}\left[b_{l,d}\right]^{2}$
@@
The complexity of obtaining variational update for $\bm{\eta}$ is then $\mathcal{O}(NL\max(D,D_{0})^{2}\max(D_{L+1},D))$.
@@
$\propto\prod_{n}^{N}\exp\left(\mathbb{E}\left[\log\frac{1}{\sqrt{|\bm{\Sigma}_{L+1}|}}\right]\right)\exp\left(\mathbb{E}\left[-\frac{1}{2}\left(\mathbf{y}_{n}-\mathbf{W}_{L+1}\mathbf{a}_{n,L}-\mathbf{b}_{L+1}\right)^{T}(\bm{\Sigma}_{L+1})^{-1}\left(\mathbf{y}_{n}-\mathbf{W}_{L+1}\mathbf{a}_{n,L}-\mathbf{b}_{L+1}\right)\right]\right)$
×$\times\prod_{l}^{L}\prod_{n}^{N}\exp\left(\mathbb{E}\left[\log\frac{1}{\sqrt{|\bm{\Sigma}_{l}|}}\right]\right)\exp\left(\mathbb{E}\left[\left(-\frac{1}{2}\left(\mathbf{a}_{n,l}-\bm{\gamma}_{n,l}\mathbf{W}_{l}\mathbf{a}_{n,l-1}-\bm{\gamma}_{n,l}\mathbf{b}_{l}\right)^{T}(\bm{\Sigma}_{l})^{-1}\left(\mathbf{a}_{n,l}-\bm{\gamma}_{n,l}\mathbf{W}_{l}\mathbf{a}_{n,l-1}-\bm{\gamma}_{n,l}\mathbf{b}_{l}\right)\right)\right]\right)$
@@
$\propto\exp\left(-\frac{1}{2}\widetilde{\mathbf{W}}_{L+1,d}\mathbf{D}^{-1}_{L+1,d}\widetilde{\mathbf{W}}_{L+1,d}^{T}-\frac{1}{2}\mathbb{E}\left[(\eta_{L+1,d})^{-2}\right]\left(\widetilde{\mathbf{W}}_{L+1,d}\left(\sum_{n}^{N}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\widetilde{\mathbf{a}}_{n,L}^{T}\right]\right)\widetilde{\mathbf{W}}_{L+1,d}^{T}-2\widetilde{\mathbf{W}}_{L+1,d}\left(\sum_{n}^{N}y_{n}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\right]\right)\right)\right)$
$\propto\exp\left(-\frac{1}{2}\left(\widetilde{\mathbf{W}}_{L+1,d}\left(\mathbf{D}^{-1}_{L+1,d}+\mathbb{E}\left[(\eta_{L+1,d})^{-2}\right]\sum_{n}^{N}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\widetilde{\mathbf{a}}_{n,L}^{T}\right]\right)\widetilde{\mathbf{W}}_{L+1,d}^{T}-2\mathbb{E}\left[(\eta_{L+1,d})^{-2}\right]\widetilde{\mathbf{W}}_{L+1,d}\left(\sum_{n}^{N}y_{n}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,L}\right]\right)\right)\right)$
@@
$\left.-\frac{1}{2}\mathbb{E}\left[(\eta_{l,d})^{-2}\right]\widetilde{\mathbf{W}}_{l,d}\left(\sum_{n}^{N}\mathbb{E}\left[\gamma_{n,l,d}^{2}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)\widetilde{\mathbf{W}}_{l,d}^{T}+\mathbb{E}\left[(\eta_{l,d})^{-2}\right]\widetilde{\mathbf{W}}_{l,d}\left(\sum_{n}^{N}\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\mathbf{a}_{n,l-1}\right]\right)\right.$
@@
$\mathbf{m}_{l,d}^{T}$ $=\mathbf{B}_{l,d}\left(\sum_{n}^{N}\left(\mathbb{E}\left[(\eta_{l,d})^{-2}\right]\mathbb{E}\left[\gamma_{n,l,d}\right]\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\widetilde{\mathbf{a}}_{n,l-1}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\left(\mathbb{E}\left[\gamma_{n,l,d}\right]-\frac{1}{2}\right)\right)\right).$
@@
The complexity of obtaining variational update for $\mathbf{W},\mathbf{b}$ is then $\mathcal{O}((L\max(D_{L+1},D)(N\max(D,D_{0})^{2}+\max(D,D_{0})^{3}))$. Assuming $\max(D,D_{0})<N$, one gets the same complexity as when updating $\bm{\eta}$, i.e. $\mathcal{O}(LN\max(D,D_{0})^{2}\max(D_{L+1},D)$.
@@
The complexity of obtaining variational update for $\bm{\omega}$ is then $\mathcal{O}((NLD\max(D,D_{0})^{2}))$.
@@
$\propto\prod_{n}^{N}\prod_{l}^{L}\prod_{l}^{D_{l}}\exp\left(-\frac{1}{2\eta_{l,d}^{2}}\mathbb{E}\left[\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\left(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d}\right)\right)^{2}\right]+\mathbb{E}\left[\frac{\gamma_{n,l,d}(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d})}{T}\right]\right).$
@@
$q(\gamma_{n,l,d})\propto$ $\exp\left(-\frac{1}{2}\mathbb{E}\left[\eta_{l,d}^{-2}\right]\left(\gamma_{n,l,d}^{2}\mathbb{E}\left[\left(\widetilde{\mathbf{W}}_{l,d}\widetilde{\mathbf{a}}_{n,l-1}\right)^{2}\right]-2\gamma_{n,l,d}\mathbb{E}\left[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\widetilde{\mathbf{W}}_{l,d}\widetilde{\mathbf{a}}_{n,l-1}\right]\right)+\frac{1}{T}\gamma_{n,l,d}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\widetilde{\mathbf{a}}_{n,l-1}\right]\right)$
$\propto$ $\exp\left(\gamma_{n,l,d}\left(-\frac{1}{2}\mathbb{E}\left[\eta_{l,d}^{-2}\right]\operatorname{Tr}\left(\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}^{T}\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)\right.\right.$
$\left.\left.+\mathbb{E}\left[\eta_{l,d}^{-2}\right]\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right)\right)$
@@
$\rho_{n,l,d}=$ $\sigma\left(\mathbb{E}\left[\eta_{l,d}^{-2}\right]\left(-\frac{1}{2}\operatorname{Tr}\left(\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}^{T}\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\widetilde{\mathbf{a}}_{n,l-1}^{T}\right]\right)+\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}\right]\right)+\frac{1}{T}\mathbb{E}\left[\widetilde{\mathbf{W}}_{l,d}\right]\mathbb{E}\left[\widetilde{\mathbf{a}}_{n,l-1}\right]\right).$
@@
The complexity of obtaining variational update for $\bm{\gamma}$ is then $\mathcal{O}(LND\max(D,D_{0})^{2})$.
@@
×$\times\prod_{n}^{N}\exp\left(-\frac{1}{2}\sum_{l}^{L}\sum_{d}^{D_{l}}\mathbb{E}\left[\frac{1}{\eta_{l,d}^{2}}\right]\mathbb{E}\left[\left({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}-\gamma_{n,l,d}\left(\mathbf{W}_{l,d}\mathbf{a}_{n,l-1}+b_{l,d}\right)\right)^{2}\right]\right)$